in a survey on study habits, students at a local college were asked whether they listened to music while studying. the following table was constructed listens to music Does not listen male 44 18 female 13 31 find the probability that a randomly selected student listens to music while studying given that the student is female.
Answer #1

probability that a randomly selected student listens to music while studying given that the student is female

=P(female and listen to music)/P(female)=13/(13+31) =13/44 =0.2955